### v1

@@ Abstract
Keywords:
Segment Anything Model Efficient Fine-tuning Semantic Segmentation
@@
Figure 2: The trade-off between MACs and segmentation accuracy (mIoU) for conventional fine-tuning methods for SAM on the ISBI2012 dataset [15]. The Red circles indicate our proposed GSAM and the triangles indicate the conventional methods. Random cropping is only performed for GSAM, cropping to the number of pixels indicated by the "size". Random cropping cannot be used except for GSAM due to its structure.
@@ 1 Introduction
Deep learning has been widely applied to various image recognition problems with great success [19, 14, 13]. In particular, in recent years, a large-scale and comprehensive model called the foundation model, has been proposed, and it is known to be a powerful model that can achieve high performance for a wide range of tasks [12, 23, 3]. In the field of semantic segmentation, SAM [18] was proposed in 2023 and can perform highly accurate segmentation on natural images without training. However, if we want to identify arbitrary classes using SAM, we need to perform fine-tuning using the teacher labels of the target dataset. Since the input image size for SAM is fixed at $1024\times 1024$, this causes a huge computational cost problem during fine-tuning.
@@
Although methods such as LoRA [16] and AdaptFormer [9] have been proposed for fine-tuning SAM more effectively, the input image size for these methods is fixed to $1024\times 1024$ is the same as for SAM, and the problem of computational cost due to the input image size has not been solved. A fine-tuning method for SAM that reduces the input image size to SAM and can train on small images such as $256\times 256$ has also been proposed [20], but again the input image size must be fixed. As the number of pixel counts varies in each dataset, the use of a fixed number of pixel counts is likely to lead to serious problems such as missing image information.
@@
In this paper, we propose Generalized SAM (GSAM), which can train even when the input image size is variable. In the conventional segmentation models based on Convolutional Neural Networks (CNN) [25, 8] which were proposed before SAM, segmentation was possible even if the input image size during training and inference were different, so it is possible to input a small random cropped image during training and input the original image size during inference to obtain segmentation results. As shown in Fig. 1, GSAM is the first method using SAM that can apply random cropping at training time, and the use of a small random cropping size reduces the computational cost at training. The fixed input size of SAM is due to fixed-size Positional Encoding. Therefore, GSAM supports variable input image sizes by employing a Positional Encoding Generator (PEG) [10] consisting of a Depth-wise Convolution layer as a substitute for Positional Encoding. Furthermore, we also propose Spatial-Multiscale (SM) AdaptFormer to consider more spatial information during fine-tuning. SM-AdaptFormer has a multi-scale structure and can handle feature vectors integrating a more diverse and wider range of spatial information. This is a segmentation-specific fine-tuning method since proper segmentation requires information at various scales.
@@
From the evaluation experiments on seven different datasets consisting of in-vehicle images, satellite images, microscopic images, endoscopic images, CT images, and transparent object images, we confirmed that the proposed GSAM can significantly reduce the computational cost of training compared to the conventional fine-tuning methods for SAM, and achieved comparable or higher segmentation accuracy. As shown in Fig. 2, GSAM achieved the trade-off of lower computational cost and higher accuracy by enabling random cropping. In particular, on the Synapse multi-organ dataset, which is CT images, GSAM achieved segmentation accuracy of more than 11% better than conventional SAM fine-tuning methods, indicating that our proposed method may be highly effective in certain areas.
@@
This paper is organized as follows. Sec. 2 describes the related works. Sec. 3 describes the details of our proposed method. Sec. 4 shows the experimental results. Finally, the conclusion and future works are described in Sec. 5.
@@ 2.1 Segmentation Models
Since U-Net [25] revolutionized the area of semantic segmentation of images, various architectures have been proposed to improve accuracy [34, 1]. Methods such as PSPNet [32] and DeepLab series [6, 7, 8], which specialize in obtaining features at various scales, and more recently, methods based on Transformer, have also emerged [5, 4]. Compared to these methods, GSAM does not require a particularly complex structure and only requires efficient fine-tuning of the foundation model, SAM, to adapt to semantic segmentation to achieve competitive performance.
@@ 2.2 Foundation Models
Since Transformer [28] was published in 2017, various foundational models have been built due to its amazing extensibility. Foundation models such as BERT [12], LLaMa [27], and GPT-4 [3] have shown ground-breaking performance in natural language processing. Recently, there has been a remarkable development of foundation models in the field of computer vision, with many high-performance models such as Segment Anything Model (SAM) [18], CLIP [23], and Stable Diffusion [24]. Among others, SAM is a segmentation model trained on huge datasets with high zero-shot generalization performance. However, foundation models generally have high generalization performance but lack expertise and require fine-tuning to recognize specific downstream tasks, and arbitrary classes properly. For this reason, a lot of research is being done to effectively and efficiently fine-tune foundation models such as SAM.
@@ 2.3 Efficient Fine-tuning for SAM
When we fine-tune foundation models with a huge number of parameters such as SAM, it is computationally very expensive to update all parameters. It is therefore common to update only a part of the weight parameters to achieve fine-tuning at a lower computational cost. Low-Rank Adaptation (LoRA) [16] successfully reduces the number of learnable parameters in downstream tasks by applying learnable low-rank matrices to each Transformer layer. This method originated in the field of natural language processing, but it has also been adapted to computer vision and can be used to fine-tune SAM. ConvLoRA [33], which applies convolutional layers to LoRA and reinforces image-related local priors to achieve higher accuracy, has also been proposed. Additionally, AdaptFormer [9] achieves higher accuracy with minimal additional learnable parameters by using two fully connected layers and an activation function in each Feed-Forward Network (FNN).
@@ 2.4 Changing The Input Image Size for SAM
Recently, some methods have been proposed that allow training with smaller images by reducing the input image size of SAM from $1024\times 1024$. SAMed [31] enables the input image size of $512\times 512$ by applying LoRA to SAM. Additionally, SAMUS [20] achieves high accuracy in medical image segmentation even with a smaller input image size of $256\times 256$ by integrating the feature maps of Transformer and CNN using Cross Attention.
@@ 3 Proposed Method
In this paper, in order to efficiently fine-tune with random cropping in training, we propose a novel method called Generalized SAM (GSAM). Fig. 3 illustrates the overview of GSAM. In GSAM, all weight parameters of the Prompt Encoder and some weight parameters of the Transformer Encoder are fixed, and the other weight parameters are updated during fine-tuning. In addition, GSAM adds a novel structure to use random cropping in training. The details of each structure of GSAM are described Sec. 3.1, Sec. 3.2.
Figure 3: Overview of Generalized SAM. $FROZEN$ indicates a network in which the weight parameters are fixed, and $Learnable$ indicates a network in which the weight parameters are updated.
@@ 3.1 Application of Random Cropping during Training
Since the input to SAM must be of a fixed size of $1024\times 1024$, it is impossible to handle random cropping with small image sizes during training. The most important reason why the input to SAM must be fixed is that the Positional Encoding in the Transformer Encoder, which is a component of SAM, is of a fixed size. Positional Encoding is a structure that adds information to each token to inform its own position of the Vision Transformer. In the case of SAM, it is a learnable weight parameter with a fixed size. Therefore, GSAM employs the Positional Encoding Generator (PEG) [10] as a substitute for Positional Encoding. PEG consists of a Depth-wise Convolution layer that considers only spatial orientation, which enables it to retain positional information even when the input size of the feature map is variable.
However, the original pretrained SAM does not support random inputs, and it is possible that global learning by Self-Attention in the Transformer Encoder alone is insufficient for small and variable inputs. Therefore, we use a new network composed of CNNs, shown in Fig. 3 as CNN Encoder in order to learn by integrating CNN features and SAM features. Since learning by local kernels of CNN is effective for smaller input images, it is considered to complement the feature map of the Transformer Encoder in SAM. GSAM adds the feature maps of the third block of ResNet101 [14], which retains some spatial information, to pre-input and post-output feature maps of the Transformer Encoder. This enables efficient fine-tuning using random cropping. From the above, the introduction of PEG and CNN Encoder enables the use of random cropping and corresponding feature extraction.
@@ 3.2 Spatial-Multiscale AdaptFormer
In order to further improve the discrimination accuracy for the target dataset, we propose Spatial-Multiscale (SM) AdaptFormer. Fig. 4 illustrates the overview of SM-AdaptFormer. AdaptFormer [9] is known as a low computational cost and high performance method for fine-tuning SAM, but AdaptFormer does not take spatial information into account. Since spatial features are important information in semantic segmentation, the proposed SM-AdaptFormer prepares multiple convolutional layers with kernels in various ranges and acquires multiscale features.
For convolutional layers with a wide range of kernels, we employ Dilated Convolution [30]. Dilated Convolution can expand the receptive field while maintaining the same kernel size, allowing global feature extraction without increasing computational cost. When Dilated Convolution is applied to the input feature map $\bm{x}$, the output feature map $\bm{y}$ can be expressed as in Eq. 1 using the position $\bm{i}$ of each pixel and the convolution kernel $\bm{w}$.
@@ 4.1 Datasets and Metrics
In the experiments, we assessed different types of image data from various domains with varying input image sizes: in-vehicle images, satellite images, microscopic images, endoscopic images, CT images, and transparent object images. We used two large datasets with more than 10,000 images and five smaller datasets with 1,000 images or less. Specifically, we used the Cityscapes dataset [11] (19 classes) for in-vehicle images and the Trans10k [29] dataset (3 classes) for transparent object images as large datasets.
As smaller datasets, we used the CamVid dataset [2] (11 classes) for in-vehicle images, the Massachusetts Buildings dataset [22] (M-Building, 2 classes) for satellite images, the ISBI2012 dataset [15] (2 classes) for microscopic images, the Kvasir-SEG dataset [17] (2 classes) for endoscopic images, and the Synapse multi-organ dataset [26] (Synapse, 9 classes) for CT images. The pixel counts for each dataset are listed in Tab. 1 and Tab. 2.
@@ 4.2 Training Conditions
In this paper, we used Pytorch library and trained the model using Adam optimizer for 200 epochs with a batch size of 8. The learning rate was initially set to 0.005 and gradually decreased using the cosine learning rate scheduler [21]. For comparison, we used conventional CNN-based networks such as U-Net [25] and DeepLabv3+ [8], as well as efficient fine-tuning methods using SAM: LoRA [16], ConvLoRA [33], AdaptFormer [9], and SAMUS [20].
For data pre-processing during training, we used random cropping, horizontal flipping, and random rotation for the CNN-based methods and GSAM. Other methods only accept images of fixed-size and therefore random cropping cannot be used. Therefore, we only applied horizontal flipping and random rotation. However, random rotation is not used for in-vehicle images and transparent object images. This is because these two types of images have a clearly defined top and bottom, and do not require pre-processing by random rotation, which would change the top and bottom direction. Image sizes for random cropping are listed in Tab. 1 and Tab. 2.
@@
U-Net [25] 67.30 72.18
DeepLabv3+ [8] 67.04 70.82
SAM [18] 57.15 83.37
LoRA [16] 59.09 85.71
ConvLoRA [33] 62.43 86.47
AdaptFormer [9] 75.49 89.91
SAMUS [20] 48.61 87.18
@@
Image size $360\times 480$ $1500\times 1500$ $256\times 256$ $224\times 224$ $224\times 224$
$\rm{Random\ crop}^{\ast}$ $256\times 256$ $512\times 512$ $128\times 128$ $224\times 224$ $224\times 224$
@@
U-Net [25] 49.79 77.73 80.01 80.99 57.31
DeepLabv3+ [8] 56.39 76.05 78.47 86.21 60.60
SAM [18] 58.27 67.59 72.15 75.94 40.61
LoRA [16] 65.20 76.76 79.18 82.20 39.08
ConvLoRA [33] 66.96 77.32 79.87 85.20 43.41
AdaptFormer [9] 74.80 80.46 80.46 88.53 61.28
SAMUS [20] 48.42 49.87 78.64 88.28 20.66
@@ 4.3 Experimental Results
Tab. 1 and Tab. 2 show the quantitative results for each dataset. Regardless of the size of the dataset, GSAM achieved comparable or even higher accuracy than existing fine-tuning methods using SAM. The red numbers in the table indicate the most accurate values. Except for the Trans10k and the CamVid datasets, the proposed methods, SM-AdaptFormer and GSAM, showed the highest accuracy for the other five datasets. Especially for the Synapse multi-organ dataset, which is CT images, SM-AdaptFormer and GSAM improved the accuracy by 4.78% and 11.50%, respectively, compared to AdaptFormer. This result indicates that our proposed method may be highly effective in certain areas. In addition, for all datasets, GSAM showed higher accuracy than the network composed of CNNs. This is considered to be due to the effectiveness of SAM itself, which is the underlying model based on Transformers, plus the learning of spatial information by the CNN-based SM-AdaptFormer and the effect of data expansion by random cropping, which took advantage of the benefits of each.
@@
Figure 5: Qualitative results. The first row is the results on the ISBI2012, the second is on the M-Building, the third is on the Cityscapes, and the fourth is on the Trans10k dataset. The Cityscapes dataset can be fed into GSAM with its original aspect ratio, but for simplicity of comparison, the same aspect ratios as other methods are shown. (a) Input image, (b) Ground truth, (c) SAM [18], (d) AdaptFormer [9], (e) SAMUS [20], (f) GSAM.
@@
Fig. 5 illustrates the qualitative results on four datasets. Specifically, the qualitative results are presented for the ISBI and the M-Building datasets, where GSAM had the highest accuracy among the comparison methods, and for the Cityscapes and the Trans10k datasets, where GSAM did not have the highest accuracy.
@@
SAM [18] 371.98 72.15
LoRA [16] 371.98 79.18
ConvLoRA [33] 511.45 79.87
AdaptFormer [9] 386.48 80.46
SAMUS [20] 145.87 78.64
GSAM(size$=256\times 256$) 270.33 80.63
GSAM(size$=128\times 128$) 74.07 80.53
GSAM(size$=64\times 64$) 18.53 78.76
GSAM(size$=32\times 32$) 7.42 71.45

### v2

@@ Abstract
Figure 2: The trade-off between MACs and segmentation accuracy (mIoU) for conventional fine-tuning methods for SAM on the ISBI2012 dataset [isbi]. The Red circles indicate our proposed GSAM and the triangles indicate the conventional methods. Random cropping is only performed for GSAM, cropping to the number of pixels indicated by the "size". Random cropping cannot be used except for GSAM due to its structure.
@@ 1 Introduction
Deep learning has been widely applied to various image recognition problems with great success [imagenet, resnet, vit]. In particular, in recent years, a large-scale and comprehensive model called the foundation model, has been proposed, and it is known to be a powerful model that can achieve high performance for a wide range of tasks [bert, clip, gpt4]. In the field of semantic segmentation, SAM [sam] was proposed in 2023 and can perform highly accurate segmentation on natural images without training. However, if we want to identify arbitrary classes using SAM, we need to perform fine-tuning using the teacher labels of the target dataset. Since the input image size for SAM is fixed at $1024\times 1024$, this causes a huge computational cost problem during fine-tuning.
@@
Although methods such as LoRA [lora] and AdaptFormer [adaptformer] have been proposed for fine-tuning SAM more effectively, the input image size for these methods is fixed to $1024\times 1024$ is the same as for SAM, and the problem of computational cost due to the input image size has not been solved. A fine-tuning method for SAM that reduces the input image size to SAM and can train on small images such as $256\times 256$ has also been proposed [samus], but again the input image size must be fixed. As the number of pixel counts varies in each dataset, the use of a fixed number of pixel counts is likely to lead to serious problems such as missing image information.
@@
In this paper, we propose Generalized SAM (GSAM), which can train even when the input image size is variable. In the conventional segmentation models based on Convolutional Neural Networks (CNN) [unet, deeplabv3plus] which were proposed before SAM, segmentation was possible even if the input image size during training and inference were different, so it is possible to input a small random cropped image during training and input the original image size during inference to obtain segmentation results. As shown in Fig.˜1, GSAM is the first method using SAM that can apply random cropping at training time, and the use of a small random cropping size reduces the computational cost at training. The fixed input size of SAM is due to fixed-size Positional Encoding. Therefore, GSAM supports variable input image sizes by employing a Positional Encoding Generator (PEG) [peg] consisting of a Depth-wise Convolution layer as a substitute for Positional Encoding. Furthermore, we also propose Spatial-Multiscale (SM) AdaptFormer to consider more spatial information during fine-tuning. SM-AdaptFormer has a multi-scale structure and can handle feature vectors integrating a more diverse and wider range of spatial information. This is a segmentation-specific fine-tuning method since proper segmentation requires information at various scales.
@@
From the evaluation experiments on seven different datasets consisting of in-vehicle images, satellite images, microscopic images, endoscopic images, CT images, and transparent object images, we confirmed that the proposed GSAM can significantly reduce the computational cost of training compared to the conventional fine-tuning methods for SAM, and achieved comparable or higher segmentation accuracy. As shown in Fig.˜2, GSAM achieved the trade-off of lower computational cost and higher accuracy by enabling random cropping. In particular, on the Synapse multi-organ dataset, which is CT images, GSAM achieved segmentation accuracy of more than 11% better than conventional SAM fine-tuning methods, indicating that our proposed method may be highly effective in certain areas.
@@
This paper is organized as follows. Sec.˜2 describes the related works. Sec.˜3 describes the details of our proposed method. Sec.˜4 shows the experimental results. Finally, the conclusion and future works are described in Sec.˜5.
@@ 2.1 Segmentation Models
Since U-Net [unet] revolutionized the area of semantic segmentation of images, various architectures have been proposed to improve accuracy [unet++, r2unet]. Methods such as PSPNet [pspnet] and DeepLab series [deeplab, deeplabv3, deeplabv3plus], which specialize in obtaining features at various scales, and more recently, methods based on Transformer, have also emerged [transunet, swinunet]. Compared to these methods, GSAM does not require a particularly complex structure and only requires efficient fine-tuning of the foundation model, SAM, to adapt to semantic segmentation to achieve competitive performance.
@@ 2.2 Foundation Models
Since Transformer [transformer] was published in 2017, various foundational models have been built due to its amazing extensibility. Foundation models such as BERT [bert], LLaMa [llama], and GPT-4 [gpt4] have shown ground-breaking performance in natural language processing. Recently, there has been a remarkable development of foundation models in the field of computer vision, with many high-performance models such as Segment Anything Model (SAM) [sam], CLIP [clip], and Stable Diffusion [sd]. Among others, SAM is a segmentation model trained on huge datasets with high zero-shot generalization performance. However, foundation models generally have high generalization performance but lack expertise and require fine-tuning to recognize specific downstream tasks, and arbitrary classes properly. For this reason, a lot of research is being done to effectively and efficiently fine-tune foundation models such as SAM.
@@ 2.3 Efficient Fine-tuning for SAM
When we fine-tune foundation models with a huge number of parameters such as SAM, it is computationally very expensive to update all parameters. It is therefore common to update only a part of the weight parameters to achieve fine-tuning at a lower computational cost. Low-Rank Adaptation (LoRA) [lora] successfully reduces the number of learnable parameters in downstream tasks by applying learnable low-rank matrices to each Transformer layer. This method originated in the field of natural language processing, but it has also been adapted to computer vision and can be used to fine-tune SAM. ConvLoRA [convlora], which applies convolutional layers to LoRA and reinforces image-related local priors to achieve higher accuracy, has also been proposed. Additionally, AdaptFormer [adaptformer] achieves higher accuracy with minimal additional learnable parameters by using two fully connected layers and an activation function in each Feed-Forward Network (FNN).
@@ 2.4 Changing The Input Image Size for SAM
Recently, some methods have been proposed that allow training with smaller images by reducing the input image size of SAM from $1024\times 1024$. SAMed [samed] enables the input image size of $512\times 512$ by applying LoRA to SAM. Additionally, SAMUS [samus] achieves high accuracy in medical image segmentation even with a smaller input image size of $256\times 256$ by integrating the feature maps of Transformer and CNN using Cross Attention.
@@ 3 Proposed Method
In this paper, in order to efficiently fine-tune with random cropping in training, we propose a novel method called Generalized SAM (GSAM). Fig.˜3 illustrates the overview of GSAM. In GSAM, all weight parameters of the Prompt Encoder and some weight parameters of the Transformer Encoder are fixed, and the other weight parameters are updated during fine-tuning. In addition, GSAM adds a novel structure to use random cropping in training. The details of each structure of GSAM are described Sec.˜3.1, Sec.˜3.2.
Figure 3: Overview of Generalized SAM. FROZENFROZEN indicates a network in which the weight parameters are fixed, and LearnableLearnable indicates a network in which the weight parameters are updated.
@@ 3.1 Application of Random Cropping during Training
Since the input to SAM must be of a fixed size of $1024\times 1024$, it is impossible to handle random cropping with small image sizes during training. The most important reason why the input to SAM must be fixed is that the Positional Encoding in the Transformer Encoder, which is a component of SAM, is of a fixed size. Positional Encoding is a structure that adds information to each token to inform its own position of the Vision Transformer. In the case of SAM, it is a learnable weight parameter with a fixed size. Therefore, GSAM employs the Positional Encoding Generator (PEG) [peg] as a substitute for Positional Encoding. PEG consists of a Depth-wise Convolution layer that considers only spatial orientation, which enables it to retain positional information even when the input size of the feature map is variable.
However, the original pretrained SAM does not support random inputs, and it is possible that global learning by Self-Attention in the Transformer Encoder alone is insufficient for small and variable inputs. Therefore, we use a new network composed of CNNs, shown in Fig.˜3 as CNN Encoder in order to learn by integrating CNN features and SAM features. Since learning by local kernels of CNN is effective for smaller input images, it is considered to complement the feature map of the Transformer Encoder in SAM. GSAM adds the feature maps of the third block of ResNet101 [resnet], which retains some spatial information, to pre-input and post-output feature maps of the Transformer Encoder. This enables efficient fine-tuning using random cropping. From the above, the introduction of PEG and CNN Encoder enables the use of random cropping and corresponding feature extraction.
@@ 3.2 Spatial-Multiscale AdaptFormer
In order to further improve the discrimination accuracy for the target dataset, we propose Spatial-Multiscale (SM) AdaptFormer. Fig.˜4 illustrates the overview of SM-AdaptFormer. AdaptFormer [adaptformer] is known as a low computational cost and high performance method for fine-tuning SAM, but AdaptFormer does not take spatial information into account. Since spatial features are important information in semantic segmentation, the proposed SM-AdaptFormer prepares multiple convolutional layers with kernels in various ranges and acquires multiscale features.
For convolutional layers with a wide range of kernels, we employ Dilated Convolution [dilatedconv]. Dilated Convolution can expand the receptive field while maintaining the same kernel size, allowing global feature extraction without increasing computational cost. When Dilated Convolution is applied to the input feature map $\bm{x}$, the output feature map $\bm{y}$ can be expressed as in Eq.˜1 using the position $\bm{i}$ of each pixel and the convolution kernel $\bm{w}$.
@@ 4.1 Datasets and Metrics
In the experiments, we assessed different types of image data from various domains with varying input image sizes: in-vehicle images, satellite images, microscopic images, endoscopic images, CT images, and transparent object images. We used two large datasets with more than 10,000 images and five smaller datasets with 1,000 images or less. Specifically, we used the Cityscapes dataset [cityscapes] (19 classes) for in-vehicle images and the Trans10k [trans10k] dataset (3 classes) for transparent object images as large datasets.
As smaller datasets, we used the CamVid dataset [camvid] (11 classes) for in-vehicle images, the Massachusetts Buildings dataset [mbuilding] (M-Building, 2 classes) for satellite images, the ISBI2012 dataset [isbi] (2 classes) for microscopic images, the Kvasir-SEG dataset [kvasir] (2 classes) for endoscopic images, and the Synapse multi-organ dataset [smo] (Synapse, 9 classes) for CT images. The pixel counts for each dataset are listed in Tab.˜1 and Tab.˜2.
@@ 4.2 Training Conditions
In this paper, we used Pytorch library and trained the model using Adam optimizer for 200 epochs with a batch size of 8. The learning rate was initially set to 0.005 and gradually decreased using the cosine learning rate scheduler [cosinelr]. For comparison, we used conventional CNN-based networks such as U-Net [unet] and DeepLabv3+ [deeplabv3plus], as well as efficient fine-tuning methods using SAM: LoRA [lora], ConvLoRA [convlora], AdaptFormer [adaptformer], and SAMUS [samus].
For data pre-processing during training, we used random cropping, horizontal flipping, and random rotation for the CNN-based methods and GSAM. Other methods only accept images of fixed-size and therefore random cropping cannot be used. Therefore, we only applied horizontal flipping and random rotation. However, random rotation is not used for in-vehicle images and transparent object images. This is because these two types of images have a clearly defined top and bottom, and do not require pre-processing by random rotation, which would change the top and bottom direction. Image sizes for random cropping are listed in Tab.˜1 and Tab.˜2.
@@
U-Net [unet] 67.30 72.18
DeepLabv3+ [deeplabv3plus] 67.04 70.82
SAM [sam] 57.15 83.37
LoRA [lora] 59.09 85.71
ConvLoRA [convlora] 62.43 86.47
AdaptFormer [adaptformer] 75.49 89.91
SAMUS [samus] 48.61 87.18
@@
Image size $360\times 480$ $1500\times 1500$ $56\times 256 22$$4\times 224 22$$24\times 224$
$\rm{Random\ crop}^{\ast}$ $256\times 256$ $512\times 512$ $128\times 128$ $\times 224 224×224$$\times 224$
@@
U-Net [unet] 49.79 77.73 80.01 80.99 57.31
DeepLabv3+ [deeplabv3plus] 56.39 76.05 78.47 86.21 60.60
SAM [sam] 58.27 67.59 72.15 75.94 40.61
LoRA [lora] 65.20 76.76 79.18 82.20 39.08
ConvLoRA [convlora] 66.96 77.32 79.87 85.20 43.41
AdaptFormer [adaptformer] 74.80 80.46 80.46 88.53 61.28
SAMUS [samus] 48.42 49.87 78.64 88.28 20.66
@@ 4.3 Experimental Results
Tab.˜1 and Tab.˜2 show the quantitative results for each dataset. Regardless of the size of the dataset, GSAM achieved comparable or even higher accuracy than existing fine-tuning methods using SAM. The red numbers in the table indicate the most accurate values. Except for the Trans10k and the CamVid datasets, the proposed methods, SM-AdaptFormer and GSAM, showed the highest accuracy for the other five datasets. Especially for the Synapse multi-organ dataset, which is CT images, SM-AdaptFormer and GSAM improved the accuracy by 4.78% and 11.50%, respectively, compared to AdaptFormer. This result indicates that our proposed method may be highly effective in certain areas. In addition, for all datasets, GSAM showed higher accuracy than the network composed of CNNs. This is considered to be due to the effectiveness of SAM itself, which is the underlying model based on Transformers, plus the learning of spatial information by the CNN-based SM-AdaptFormer and the effect of data expansion by random cropping, which took advantage of the benefits of each.
@@
Figure 5: Qualitative results. The first row is the results on the ISBI2012, the second is on the M-Building, the third is on the Cityscapes, and the fourth is on the Trans10k dataset. The Cityscapes dataset can be fed into GSAM with its original aspect ratio, but for simplicity of comparison, the same aspect ratios as other methods are shown. (a) Input image, (b) Ground truth, (c) SAM [sam], (d) AdaptFormer [adaptformer], (e) SAMUS [samus], (f) GSAM.
@@
Fig.˜5 illustrates the qualitative results on four datasets. Specifically, the qualitative results are presented for the ISBI and the M-Building datasets, where GSAM had the highest accuracy among the comparison methods, and for the Cityscapes and the Trans10k datasets, where GSAM did not have the highest accuracy.
@@
SAM [sam] 371.98 72.15
LoRA [lora] 371.98 79.18
ConvLoRA [convlora] 511.45 79.87
AdaptFormer [adaptformer] 386.48 80.46
SAMUS [samus] 145.87 78.64
[…]